Seven inches are equivalent to 17.78 cm. Find the equation converting length in cm to length in inches. What is the length in inches of a 96.2 cm bar?
Answer by checkley71(8403)

7/17.78=x/96.2 or 17.78x=673.4 or x=673.4/17.78 or x=37.874in